# Explainable Concept Generation through Vision-Language Preference Learning

### v1

@@ 4.3 Are the generated concepts reliable?
Figure 8: C-deletion. Removing concepts over time to measure the reliability. Colored values indicate the cumulative avg. accuracy over time (the lower the better as it indicates important concepts are removed sooner)
After finding the relationship between generated concepts and input images, it is important to validate the importance of the identified concepts. To do that, we applied c-deletion to the class images for each identified concept. We tested it with, 1) traditional c-deletion, where we gradually delete concept segments based on decreasing order of TCAV importance and 2) on a modified c-deletion, which stops deleting after removing the concept we test. The results for the c-deletion, original and clipped are shown in the Fig. 8.
4.4 Applications and generalizability
We discussed how RLPO can be used as a diagnostic tool for the engineers. As a specific application, we see what concepts are removed and added as well as how the concept importance changes when we fine-tune ResNet50 on ImageNet to improve accuracy (Fig. 6).
5 Limitations and Conclusions
The process of navigating an infinitely large concept space and generating explainable concepts from textual inputs presents several challenges, particularly when dealing with complex, high-level concepts. We showed how DRL and preference optimization can be combined to efficiently navigate this space. However, RLPO also suffers from several limitations. First, this analysis cannot be performed real-time since generating images from SD, learning the DQN, and fine-tuning the SD with preferences takes some time. Also, the concepts that our algorithm generates can be diverse as it tries to reveal the concepts inherent to the NUT, making it less domain-specific (e.g., for a medical application, there is a chance it might generate non-medical images if the NUT activations get excited for non-medical data). As a future extension, we hope to input preferences from both TCAV and application experts while optimizing, making generated explanations even more aligned to specific applications. Despite the challenges, our results show how to leverage the strengths of visual representations and adaptive learning to provide intuitive and effective solutions for understanding complex, high-level concepts in black-box neural networks.
Broader Imapcts: Our paper is a generic algorithmic contribution and we do not foresee direct negative societal impacts. As a positive trait, the model can potentially be used to explain algorithmic bias.

### v2

@@ D.3.4 C-deletion
Figure 25: C-deletion. Removing concepts over time to measure the reliability. The colored numbers indicate the area under the curve (the lower the better).
C-deletion evaluations assesses the impact of removing certain contextual inputs (features, variables, or states) on a model’s performance as shown in Fig. 25.
D.4 RLPO in sentiment analysis
We extended our method to the NLP domain, successfully identifying which parts of the input contribute to specific outputs. For sentiment analysis, a binary classification problem, we present results for both positive and negative classes.
A list of positive and negative prompts was created, analogous to class images in traditional image classification tasks. Random prompts, similar to those in Fig. 12, were used to simulate random classes. Every word in the prompt, excluding stop words, along with its synonyms, was treated as a concept for this experiment. Synonyms were generated using the Mistral-7B Instruct model, serving a role comparable to the image generation model in image-based settings. We observed that multiple words were identified, along with their influence on the overall prompt, for both classes as shown in Fig. 10 and Fig. 26.
Figure 26: Generated concepts explain why a given text is classified as a negetive sentiment.
D.5 Effects of Fine-tuning
When fine-tuning a model, the optimization process updates its weights through gradient-based methods, causing shifts in the concepts (Fig. 9) it learns. These weight adjustments modify how the model attends to different regions or patterns in an image, leading to changes in the internal activation maps and the conceptual understanding of the input. As the model learns new concepts or refines existing ones, it adjusts its feature extraction and decision-making processes to better align with the specific objectives of the fine-tuning task, thereby altering the way it interprets and generates outputs.
To demonstrate this experiment, a ResNet50 model, defined in Appendix C.1, was used. We fine-tuned only the final layer of the neural network by setting the learning rate to 0.001 and momentum to 0.9 using the Stochastic Gradient Descent (SGD) optimizer. Maintaining a low learning rate was crucial to preserving high accuracy. RLPO was then applied to the fine-tuned ResNet50, and a shift in the learned concepts was observed, as shown in Fig. 9. This process highlights the model’s sensitivity to fine-tuning and how training on a subset can shift its conceptual interpretation of images.
D.6 Human survey
The survey involved 50 participants, each of whom was shown 10 class images along with two concept options as shown in Fig. 27: one derived from a retrieval-based method and the other generated using RLPO. The participants were divided into Laymen and Experts.
  1. 1.
    Expert: Computer science graduates who are familiar with the concept of explainability and have a working knowledge of AI or machine learning systems.
  2. 2.
    Laymen: Individuals without expertise in computer science, AI, or explainability, representing the general public’s perspective.
Figure 27: A screenshot from our human survey with instructions and a sample question.